Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

A pumping station at a hydroelectric plant operates two identical pumps. Each pump has a probability of failure of 0.06, and the probability that both pumps fail is 0.0036. (a) Are failures in the two pumps mutually exclusive? b) What is the probability that at least one of the pumps fails? (c) Are failures in the two pumps independent?

OpenStudy (anonymous):

a. P(A and B) = 0.0036 Check the conditional probability: P(A|B) = P(A and B)/P(B) (0.0036)/(0.06) does not equal P(A and B). Therefore, they are not mutually exclusive. b. P(At least one fails) = P(A or B fails) P(A or B failing) = P(A) + P(B) - P(A and B) = (0.06) + (0.06) - (0.06*0.06) c. Yes, they are independent of each other because (0.06)(0.06) = 0.0036, which is the same as P(A and B), and is the definition of independent events.

OpenStudy (anonymous):

[a] No [b] Probability of all not failing = 0.94*0.94=(1-0.06)^2 = 0.8836 Probability of at least one fails = 1-0.8836 = 0.1164 [c] Yes

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!